# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

### v2

@@ 4.1 Baselines Selection
Experiment Results. As shown in Table 5, in the long-tailed semi-supervised setting, SSL methods are designed to ensure that the learning process effectively captures all classes, even when the distribution is dominated by a few head classes. However, as shown in Table 1, their performance often declines in the same datasets compared to standard SSL, highlighting their ongoing vulnerability to the challenges of long-tailed distributions. In contrast, the evaluation results of pre-trained VLMs underscore their efficiency, which can be fine-tuned with only a small amount of labeled data and still perform well, even on low-resolution datasets, a common challenge for these models. As shown in Table 6, pre-trained models perform well on highly imbalanced datasets like STL-10 but struggle on semantically coarse datasets such as ImageNet-127 cossl . ImageNet-127 is obtained by collapsing the 1,000 fine-grained ImageNet-1K classes into 127 coarse-grained categories according to their positions in the WordNet semantic hierarchy, yielding a long-tailed distribution (e.g., grouping “abacus,” “acoustic guitar,” and “analog clock” under the broader category “device”). Zero-shot CLIP evaluations reveal that this hierarchical coarsening injects semantic ambiguity, making it challenging for pretrained VLMs to confidently associate a fine-grained concept with a specific image. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. Detailed analysis and visualization are shown in Appendix C.
@@ 5 Discussion
Potential for combining SSL and pre-trained models. Integrating SSL with pre-trained models offers new opportunities for semi-supervised classification. While SSL leverages unlabeled data to reduce labeling costs, pre-trained models excel in zero-shot learning and generalization. Their combination can effectively address challenges such as few-shot learning, class imbalance, and other specialized scenarios. For instance, PromptKD promptkd introduces a two-stage, prompt-driven knowledge distillation framework that aligns teacher and student models through learnable prompts, achieving effective knowledge transfer using a diverse mix of ID and OOD unlabeled images, resulting in significant improvements across various datasets. Similarly, UPL huang2022unsupervised automatically generates high-quality prompts from unlabeled data, reducing the dependence on large-scale labeled datasets and improving overall model fit. GRIP menghini2023enhancing represents another promising direction, proposing a pseudo-label-based fine-tuning method that effectively combines SSL principles with VLMs. Clearly, leveraging pseudo-labeled data, semi-supervised learning concepts, and the generalization strengths of VLMs offers a promising path for enhancing performance across a wide range of downstream tasks, prompting a fresh perspective on hybrid learning strategies.
Effects of training with long-tailed data on pre-trained models. When training data follows a long-tailed distribution, excessive fine-tuning of pre-trained models can be detrimental, particularly to the performance on tail classes. As highlighted in lift , aggressive fine-tuning can disrupt the class-conditional structure learned during pre-training, resulting in poor representation for underrepresented classes. To address this issue, they proposed LIFT, a lightweight fine-tuning approach that modifies only a small portion of the model (e.g., prompts or adapter layers), effectively adapting to long-tailed distributions while preserving the semantic consistency of the pre-trained model. This demonstrates that leveraging pre-trained models can help mitigate the adverse effects of long-tailed distributions. However, current approaches primarily focus on reducing fine-tuning intensity, often at the cost of head class performance. This highlights the need for strategies that can guide models to fully adapt to imbalanced data while minimizing learning bias.

### v3

@@ 4.1 Baselines Selection
Experiment Results. As shown in Table 5, in the long-tailed semi-supervised setting, SSL methods are designed to ensure that the learning process effectively captures all classes, even when the distribution is dominated by a few head classes. However, as shown in Table 1, their performance often declines in the same datasets compared to standard SSL, highlighting their ongoing vulnerability to the challenges of long-tailed distributions. In contrast, the evaluation results of pre-trained VLMs underscore their efficiency, which can be fine-tuned with only a small amount of labeled data and still perform well, even on low-resolution datasets, a common challenge for these models. As shown in Table 6, pre-trained models perform well on highly imbalanced datasets like STL-10 but struggle on semantically coarse datasets such as ImageNet-127 [12]. ImageNet-127 is obtained by collapsing the 1,000 fine-grained ImageNet-1K classes into 127 coarse-grained categories according to their positions in the WordNet semantic hierarchy, yielding a long-tailed distribution (e.g., grouping “abacus,” “acoustic guitar,” and “analog clock” under the broader category “device”). Zero-shot CLIP evaluations reveal that this hierarchical coarsening injects semantic ambiguity, making it challenging for pretrained VLMs to confidently associate a fine-grained concept with a specific image. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. Detailed analysis and visualization are shown in Appendix C.
@@ 5 Discussion
Potential for combining SSL and pre-trained models. Integrating SSL with pre-trained models offers new opportunities for semi-supervised classification. While SSL leverages unlabeled data to reduce labeling costs, pre-trained models excel in zero-shot learning and generalization. Their combination can effectively address challenges such as few-shot learning, class imbalance, and other specialized scenarios. For instance, PromptKD [34] introduces a two-stage, prompt-driven knowledge distillation framework that aligns teacher and student models through learnable prompts, achieving effective knowledge transfer using a diverse mix of ID and OOD unlabeled images, resulting in significant improvements across various datasets. Similarly, UPL [23] automatically generates high-quality prompts from unlabeled data, reducing the dependence on large-scale labeled datasets and improving overall model fit. GRIP [36] represents another promising direction, proposing a pseudo-label-based fine-tuning method that effectively combines SSL principles with VLMs. Clearly, leveraging pseudo-labeled data, semi-supervised learning concepts, and the generalization strengths of VLMs offers a promising path for enhancing performance across a wide range of downstream tasks, prompting a fresh perspective on hybrid learning strategies.
Effects of training with long-tailed data on pre-trained models. When training data follows a long-tailed distribution, excessive fine-tuning of pre-trained models can be detrimental, particularly to the performance on tail classes. As highlighted in [47], aggressive fine-tuning can disrupt the class-conditional structure learned during pre-training, resulting in poor representation for underrepresented classes. To address this issue, they proposed LIFT, a lightweight fine-tuning approach that modifies only a small portion of the model (e.g., prompts or adapter layers), effectively adapting to long-tailed distributions while preserving the semantic consistency of the pre-trained model. This demonstrates that leveraging pre-trained models can help mitigate the adverse effects of long-tailed distributions. However, current approaches primarily focus on reducing fine-tuning intensity, often at the cost of head class performance. This highlights the need for strategies that can guide models to fully adapt to imbalanced data while minimizing learning bias.